# Exploiting Symmetry in Dynamics for Model-Based Reinforcement Learning with Asymmetric Rewards

### v1

@@ Abstract
Recent work in reinforcement learning has leveraged symmetries in the model to improve sample efficiency in training a policy. A commonly used simplifying assumption is that the dynamics and reward both exhibit the same symmetry. However, in many real-world environments, the dynamical model exhibits symmetry independent of the reward model: the reward may not satisfy the same symmetries as the dynamics. In this paper, we investigate scenarios where only the dynamics are assumed to exhibit symmetry, extending the scope of problems in reinforcement learning and learning in control theory where symmetry techniques can be applied. We use Cartan’s moving frame method to introduce a technique for learning dynamics which, by construction, exhibit specified symmetries. We demonstrate through numerical experiments that the proposed method learns a more accurate dynamical model.
@@ I Introduction
A significant area of research in reinforcement learning (RL) is in improving sample efficiency, the amount a learning method must interact with the environment to learn a good policy. While model-free methods such as [lillicrap2019continuous], [haarnoja_soft_2018], and [schulman2017proximal] have shown the ability to achieve high rewards in many complex environments, they typically require many environment interactions. Model-based reinforcement learning (MBRL) methods, on the other hand, have shown promise in learning policies quickly relative to the number of environment interactions. This improvement in sample efficiency confirms that models contain a great deal of information useful for improving policies [moerland_mbrlsurvey].
@@
A common assumption in the use of symmetry in both RL and optimal control is that both the dynamics and reward (or cost) function exhibit the same symmetries. This combination of assumptions results in the optimal policy (or control law) likewise being symmetric [ravindran_model_2002, maidens_dynamic_programming]. The symmetric policy structure is exploited for Markov Decision Processes (MDPs) and RL in [vanderpol2021mdp, wang_mathrmso2-equivariant_2022, ravindran_model_2002], [zinkevich_symmetry_2001], and for dynamic programming in [maidens_dynamic_programming].
In many scenarios, however, symmetries in the dynamical model are not necessarily applicable to the reward model. For example, multiple tasks might be executed in the same environment, where the same dynamical symmetries exist, while the reward model changes from task to task. Exploiting symmetries only in the dynamical model widens the scope of RL problems to which symmetry can be applied.
@@
This paper is organized as follows: in Section II, we provide background information on symmetries, and then use Cartan’s method of moving frames to relate a symmetrical dynamical model to a function operating on an input space of reduced dimension. In Section III, we use the result of Section II to learn a dynamical model which, by construction, satisfies specified symmetries. In Section IV we apply symmetry reduction to learn dynamical models for two examples having symmetry in dynamics.
@@ II Symmetries in Model Dynamics
where $x_{k}\in\mathcal{X}=\mathbb{R}^{n}$ is the state, $u_{k}\in\mathcal{U}=\mathbb{R}^{n_{u}}$ is the control input, and $k\in\mathbb{N}$ denotes the time step.
Following [SymmetryPreservingObservers] and [maidens_reachability], we provide definitions of a transformation group and of dynamics invariant under a transformation group.
Definition 1 (Transformation Group)
Let $G$ be a group. A transformation group on $\mathcal{X}\times\mathcal{U}$ is a set of tuples $(\phi_{\alpha},\psi_{\alpha})$ parameterized by $\alpha\in G$ such that $\phi_{\alpha}:\mathcal{X}\to\mathcal{X}$ and $\psi_{\alpha}:\mathcal{U}\to\mathcal{U}$ are diffeomorphisms that satisfy the following criteria for all $x\in\mathcal{X}$, $u\in\mathcal{U}$, and $g_{1},g_{2}\in G$:
  • $\phi_{e}(x)=x$ and $\psi_{e}(u)=u$ where $e$ is the identity of group $G$ and
@@
  • $\phi_{g_{1}}(\phi_{g_{2}}(x))=\phi_{g_{1}\cdot g_{2}}(x)$ and $\psi_{g_{1}}(\psi_{g_{2}}(u))=\psi_{g_{1}\cdot g_{2}}(u)$ where $g_{1}\cdot g_{2}$ denotes the group operation. We will denote this simply by $g_{1}g_{2}$ henceforth.
Definition 2
We say the system (1) is invariant to the transformation group $\{(\phi_{\alpha},\psi_{\alpha})\}_{\alpha\in G}$, or $G$-invariant, if
$F(\phi_{\alpha}(x),\psi_{\alpha}(u))=\phi_{\alpha}(F(x,u))$ (2)
for all $x\in\mathcal{X}$, $u\in\mathcal{U}$, and $\alpha\in G$.
We now use Cartan’s Moving Frame method [cartan1937théorie] to parameterize the system (1) in terms of a function with lower dimensional input space, when (1) is invariant to a transformation group parameterized by a Lie group.
Let $G$ be a Lie group which acts on the state space with diffeomorphisms $\{\phi_{\alpha}\}_{\alpha\in G}$, and acts on the control input space with diffeomorphisms $\{\psi_{\alpha}\}_{\alpha\in G}$. Assume that $G$ is $r$-dimensional, with $r\leq n$. Now, split $\phi_{\alpha}$ into $(\phi_{\alpha}^{a},\phi_{\alpha}^{b})$ with $r$ and $n-r$ components respectively, such that $\phi_{\alpha}^{a}$ is invertible with respect to $\alpha$. Pick $c$ that is in the range of $\phi^{a}$, and define $\mathcal{C}=\left\{x|\phi_{e}^{a}(x)=c\right\}$ where $e\in G$ is the group identity. Note that $\mathcal{C}$ is $n-r$ dimensional. This is the lower dimensional space that will be used to parameterize system (1).
Assume that, for all $x$, there exists a unique $\alpha\in G$ such that $\phi_{\alpha}(x)\in\mathcal{C}$. Then, we define a function $\gamma$ such that $\phi_{\gamma(x)}(x)\in\mathcal{C}$. Using the function $\gamma$, which is called a moving frame, we define the following map, $\rho:\mathcal{X}\to\mathcal{X}^{b}$:
@@
Let $\alpha\in G$. Then,
$\rho(\phi_{\alpha}(x))$ $=\phi_{\gamma(\phi_{\alpha}(x))}^{b}(\phi_{\alpha}(x))$
$=\phi_{\gamma(\phi_{\alpha}(x))\alpha}^{b}(x).$
From the definition of $\gamma$, we have that $\phi_{\gamma(x)}^{a}(x)=c$. Note that $\phi_{\gamma(\phi_{\alpha}(x))}^{a}(\phi_{\alpha}(x))=\phi_{\gamma(\phi_{\alpha}(x))\alpha}^{a}(x)$. Therefore, $\gamma(\phi_{\alpha}(x))\alpha=\gamma(x)$. Plugging this into the expression for $\rho(\phi_{\alpha}(x))$ gives $\rho(\phi_{\alpha}(x))=\phi_{\gamma(x)}^{b}(x)=\rho(x)$. ∎
The map $\rho$ outputs the element of $\mathcal{C}$, eliminating the part equal to $c$, that is related by a symmetry transformation to the state $x$. This effectively transforms the state space from dimension $n$ to dimension $n-r$. The following theorem shows the dynamics can be evaluated using the lower dimensional output of $\rho$.
@@
The system (1), $F:\mathcal{X}\times\mathcal{U}\to\mathcal{X}$, is $G$-invariant if and only if there exists a map $\bar{F}:\mathcal{X}^{b}\times\mathcal{U}\to\mathcal{X}$ such that
@@
Assume that the dynamics $F$ are $G$-invariant. Note that $\rho$ restricted to $\mathcal{C}$ is invertible. Define $\bar{F}$ by $\bar{F}(\bar{x},u)=F(\rho|_{\mathcal{C}}^{-1}(\bar{x}),u)$. Letting $x=\rho|_{\mathcal{C}}^{-1}(\bar{x})$, this can equivalently be written as $\bar{F}(\rho(x),u)=F(x,u)$ when $x\in\mathcal{C}$. Since $F$ is $G$-invariant, $\phi_{\gamma(x)}(F(x,u))=F(\phi_{\gamma(x)}(x),\psi_{\gamma(x)}(u))$. Note that $\phi_{\gamma(x)}(x)\in\mathcal{C}$. This can be rewritten in terms of $\bar{F}$ as $\bar{F}(\rho(\phi_{\gamma(x)}(x)),\psi_{\gamma(x)}(u))$. Since $\rho$ is invariant to action of $G$, this equals $\bar{F}(\rho(x),\psi_{\gamma(x)}(u))$, as desired.
Now we prove the reverse direction. Assume there exists $\bar{F}:\mathcal{X}^{b}\times\mathcal{U}\to\mathcal{X}$ such that (4) for all $x\in\mathcal{X}$, $u\in\mathcal{U}$. Let $\alpha\in G$. Then,
$F($ $\phi_{\alpha}(x),\psi_{\alpha}(u))$
$=\phi_{\gamma(\phi_{\alpha}(x))}^{-1}(\bar{F}(\rho(\phi_{\alpha}(x)),\psi_{\gamma(\phi_{\alpha}(x))}(\psi_{\alpha}(u))))$
$=\phi_{\gamma(\phi_{\alpha}(x))}^{-1}(\bar{F}(\rho(x),\psi_{\gamma(\phi_{\alpha}(x))\alpha}(u))).$
From the proof of Lemma 1, we know $\gamma(\phi_{\alpha}(x))\alpha=\gamma(x)$. Therefore the above can be simplified to
$F(\phi_{\alpha}(x),\psi_{\alpha}(u))$ $=\phi_{\gamma(x)\alpha^{-1}}^{-1}(\bar{F}(\rho(x),\psi_{\gamma(x)}(u)))$
$=\phi_{\alpha}(\phi_{\gamma(x)^{-1}}(\phi_{\gamma(x)}(F(x,u))))$
$=\phi_{\alpha}(F(x,u)).$
@@ III Learning Model Dynamics
A critical part of the theory in Section II is that it depends only on the state space, control input space, and the symmetries of the system; we do not require knowledge of the system dynamics.
We learn a dynamical model that is invariant to a given transformation group with the help of Theorem 1. Unlike a standard formulation for learning a dynamical model, which would learn $F:\mathcal{X}\times\mathcal{U}\to\mathcal{X}$, we instead learn the function $\bar{F}:\mathcal{X}^{b}\times\mathcal{U}\to\mathcal{X}$ from the statement of Theorem 1. With this function $\bar{F}$, we construct the model dynamics $F$ as
@@
The relationship between $F$ and $\bar{F}$ is depicted in Figure 1. From Theorem 1, $F$ is, by construction, $G$-invariant. A side benefit is that training is done with a lower dimensional input space ($\mathcal{X}^{b}\times\mathcal{U}$ instead of $\mathcal{X}\times\mathcal{U}$). To learn this function $\bar{F}$, given a dataset $\mathcal{D}$ consisting of tuples $(x_{k},u_{k},x_{k+1})$, we train $\bar{F}$ to map $(\rho(x_{k}),\psi_{\gamma(x_{k})}(u_{k}))$ to $\phi_{\gamma(x_{k})}(x_{k+1})$.
@@
Note in practice it is typical to learn a function which maps $(x_{k},u_{k})$ to $x_{k+1}-x_{k}$ instead of $x_{k+1}$ directly. To do this, we learn $\Delta\bar{F}$ to map $(\rho(x_{k}),\psi_{\gamma(x_{k})}(u_{k}))$ to $\phi_{\gamma(x_{k})}(x_{k+1})-\phi_{\gamma(x_{k})}(x_{k})$. Then, we construct $F$ by
@@
This construction for $F$ is $G$-invariant since it is equivalent to using $\bar{F}(\bar{x},\bar{u})\triangleq\Delta\bar{F}(\bar{x},\bar{u})+\rho|_{\mathcal{C}}^{-1}(\bar{x})$, which satisfies the condition in Theorem 1. Finally, $\Delta F(x,u)$ can defined as $\Delta F(x,u)\triangleq F(x,u)-x$ to provide the same $x_{k+1}-x_{k}$ interface expected by many libraries. This construction for $\Delta F$ in terms of $\Delta\bar{F}$ expands as follows.
@@
When $\phi_{\alpha}$ is a homomorphism with respect to $+$ for all $\alpha\in G$, (6) simplifies to
@@
An example where this occurs is when $\phi_{\alpha}$ is a linear transformation.

### v2

@@ Abstract
Recent work in reinforcement learning has leveraged symmetries in the model to improve sample efficiency in training a policy. A commonly used simplifying assumption is that the dynamics and reward both exhibit the same symmetry; however, in many real-world environments, the dynamical model exhibits symmetry independent of the reward model. In this paper we assume only the dynamics exhibit symmetry, extending the scope of problems in reinforcement learning and learning in control theory where symmetry techniques can be applied. We use Cartan’s moving frame method to introduce a technique for learning dynamics which, by construction, exhibit specified symmetries. Numerical experiments demonstrate that the proposed method learns a more accurate dynamical model.
@@ I Introduction
A major research area in reinforcement learning (RL) is improving sample efficiency, the amount of interaction with the environment needed to learn a good policy. While model-free methods such as [lillicrap2019continuous, haarnoja_soft_2018, schulman2017proximal] have shown the ability to achieve high rewards in many complex environments, they require many environment interactions. Model-based reinforcement learning (MBRL) methods, on the other hand, have shown promise in learning policies quickly relative to the number of environment interactions. This improvement in sample efficiency confirms that models contain a great deal of information useful for improving policies [moerland_mbrlsurvey].
@@
A common assumption when using symmetry in RL and optimal control is that both the dynamics and reward (or cost) function exhibit the same symmetries, which results in the optimal policy (or control law) likewise being symmetric [ravindran_model_2002, maidens_dynamic_programming]. The symmetric policy structure is exploited for Markov Decision Processes (MDPs) and RL in [vanderpol2021mdp, wang_mathrmso2-equivariant_2022, ravindran_model_2002], [zinkevich_symmetry_2001], and for dynamic programming in [maidens_dynamic_programming]. In many scenarios, however, symmetries in the dynamical model are not necessarily applicable to the reward model. For example, multiple tasks might be executed in the same environment, where the same dynamical symmetries exist, while the reward model changes from task to task. Exploiting symmetries only in the dynamical model widens the scope of RL problems to which symmetry can be applied.
@@
This paper is organized as follows: in Section II, we provide background information on symmetries, and then use Cartan’s method of moving frames to relate a symmetrical dynamical model to a function operating on an input space of reduced dimension. In Section III, we use the result of Section II to learn a dynamical model which, by construction, satisfies specified symmetries. In Section IV we apply symmetry reduction to learn dynamical models for two examples.
@@ II Symmetries in Model Dynamics
where $x_{k}\in\mathcal{X}=\mathbb{R}^{n}$ is the state, $u_{k}\in\mathcal{U}=\mathbb{R}^{n_{u}}$ is the control input, and $k\in\mathbb{N}$ denotes the time step. Symmetry refers to a structure of the dynamical system such that transformations of the input $(x_{k},u_{k})$ to $F$ result in corresponding, known, transformations to the output $x_{k+1}$. We formalize this notion with Lie transformation groups, an introduction to which can be found in [bulloGeometricControlMechanical2004] and [leeIntroductionSmoothManifolds2012].
Definition 1 (Lie Transformation Group)
Let $G$ be a Lie group acting on a smooth manifold $\Sigma$. Denote the group action by $w(g,\sigma)\mapsto g\cdot\sigma$. The group $G$ is a Lie transformation group if $w:G\times\Sigma\to\Sigma$ is smooth.
Important properties of the group action are:
  • $e\cdot\sigma=\sigma$ for all $\sigma$ where $e$ is the identity of group $G$.
@@
  • $g_{1}\cdot(g_{2}\cdot\sigma)=(g_{1}*g_{2})\cdot\sigma$ for all $g_{1},g_{2}\in G,\sigma\in\Sigma$, where $g_{1}*g_{2}$ denotes the group operation, which we will denote by $g_{1}g_{2}$ henceforth.
We represent symmetries of (1) with a $r$-dimensional Lie group $G$ which is a Lie transformation group of both $\mathcal{X}$ and $\mathcal{U}$. Assume $r\leq n$. For notation, we consider the maps $\phi_{g}:\mathcal{X}\to\mathcal{X}$ defined by $x\mapsto g\cdot x$ and $\psi_{g}:\mathcal{U}\to\mathcal{U}$ defined by $u\mapsto g\cdot u$. Since the group actions are smooth, both of these maps are smooth for all $g\in G$. Further, they have smooth inverses $\phi^{-1}_{g}=\phi_{g^{-1}}$ and $\psi^{-1}_{g}=\psi_{g^{-1}}$.
We next define dynamics invariant to symmetry:
Definition 2 (Invariant Dynamics)
The system (1) is $G$-invariant if $F(\phi_{g}(x),\psi_{g}(u))=\phi_{g}(F(x,u))$ for all $g\in G$, $x\in\mathcal{X}$, $u\in\mathcal{U}$.
The premise of our method is to transform $(x,u)$ to its canonical form, corresponding to an element of $\{(\phi_{g}(x),\psi_{g}(u))\ |\ g\in G\}$. When $F$ in (1) is $G$-invariant, we parameterize it by a function that operates on these canonical forms, which reside in a lower-dimensional space.
II-A Cartan’s Moving Frame
We use Cartan’s moving frames [cartan1937théorie] to construct a map from elements of the state space to their canonical forms. What follows is a brief recap of Cartan’s method based on [SymmetryPreservingObservers, maidens_reachability, olverClassicalInvariantTheory1999]. Note that, in general, this method constructs a map which exists only locally, due to the use of the implicit function theorem. On many practical examples, such as those in this work and [maidens_reachability], the map exists over all $\mathcal{X}$ except a lower-dimensional submanifold. Thus, in our presentation, we will focus on transformations within a single smooth chart. Refer to [olverClassicalInvariantTheory1999] for a thorough treatment of the subject.
Consider the Lie group $G$ and its action on $\mathcal{X}$ as before. Assume $\mathcal{X}$ can be split into $\mathcal{X}^{a}$ and $\mathcal{X}^{b}$ with $r$ and $n-r$ dimensions respectively such that, for some $x_{0}\in\mathcal{X}$, the map from $g\in G$ to the projection of $g\cdot x_{0}$ onto $\mathcal{X}^{a}$ is invertible. Pick $c\in\mathcal{X}^{a}$ in the range of this map. For all $g\in G$, let $\phi_{g}^{a}:\mathcal{X}\to\mathcal{X}^{a}$ and $\phi_{g}^{b}:\mathcal{X}\to\mathcal{X}^{b}$ denote the projections of $\phi_{g}$ onto $\mathcal{X}^{a}$ and $\mathcal{X}^{b}$ respectively. Assume that, for all $x\in\mathcal{X}$, there exists a unique $g\in G$ such that $\phi_{g}^{a}(x)=c$. The set $\mathcal{C}=\{x|\phi_{e}^{a}(x)=c\}$ is an $n-r$ dimensional submanifold of $\mathcal{X}$ called a cross-section. We can then define the unique map $\gamma:\mathcal{X}\to G$, such that $\phi_{\gamma(x)}(x)\in\mathcal{C}$. This $\gamma$ is called a moving frame. Using $\gamma$, we define the map $\rho:\mathcal{X}\to\mathcal{X}^{b}$ from elements in $\mathcal{X}$ to their canonical forms:
@@ II-A Cartan’s Moving Frame
Let $g\in G$. Then, $\rho(\phi_{g}(x))=\phi_{\gamma(\phi_{g}(x))}^{b}(\phi_{g}(x))=\phi_{\gamma(\phi_{g}(x))g}^{b}(x)$. From the definition of $\gamma$, we have that both $\phi_{\gamma(x)}^{a}(x)$ and $\phi_{\gamma(\phi_{g}(x))}^{a}(\phi_{g}(x))$ equal $c$. Note that $\phi_{\gamma(\phi_{g}(x))}^{a}(\phi_{g}(x))=\phi_{\gamma(\phi_{g}(x))g}^{a}(x)$. Since $\gamma(x)$ is the unique group element $h$ such that $\phi_{h}^{a}(x)=c$, we have that $\gamma(\phi_{g}(x))g=\gamma(x)$. Plugging this into the expression for $\rho(\phi_{g}(x))$ gives $\rho(\phi_{g}(x))=\phi_{\gamma(x)}^{b}(x)=\rho(x)$. ∎
Example 1
In this example, we demonstrate solving for the moving frame for a system with both translation and rotation symmetries. Consider a car with state $x=(y,z,v_{y},v_{z},h_{y},h_{z})$ where $y$ and $z$ correspond to positions in a plane, $v_{y}$ and $v_{z}$ correspond to velocities along the $y$ and $z$ directions, and $h_{y}$ and $h_{z}$ correspond to the cosine and sine of the heading angle (and thus $h_{y}^{2}+h_{z}^{2}=1$). Since we assume the dynamics of the car are both translation- and rotation-invariant, we consider the special Euclidean group $G=\mathrm{SE}(2)$, and parameterize it by a translation $(y^{\prime},z^{\prime})$ and a rotation $\theta^{\prime}$. Then, with rotation matrix $R_{\theta^{\prime}}=\begin{bmatrix}\cos\theta^{\prime}&-\sin\theta^{\prime}\\ \sin\theta^{\prime}&\cos\theta^{\prime}\end{bmatrix}$, we define the action of $G$ on the state space by
$\phi_{(y^{\prime},z^{\prime},\theta^{\prime})}(x)=\begin{bmatrix}R_{\theta^{\prime}}&0&0\\ 0&R_{\theta^{\prime}}&0\\ 0&0&R_{\theta^{\prime}}\end{bmatrix}x+\begin{bmatrix}\begin{bmatrix}y^{\prime}\\ z^{\prime}\end{bmatrix}\\ 0\\ 0\end{bmatrix}.$ (3)
We now derive $\gamma$ and $\rho$. Let $\mathcal{X}^{a}$ be the projection of $\mathcal{X}$ onto the $(y,z,h_{y},h_{z})$ components. Pick $c=(0,0,1,0)$, representing a position at the origin with a heading angle of 00. There exists a unique $(y^{\prime},z^{\prime},\theta^{\prime})$ to transform each $x\in\mathcal{X}$ into the cross-section defined by $\mathcal{C}$. We solve the following system of equations to find the moving frame:
$\begin{bmatrix}0\\ 0\\ 1\\ 0\end{bmatrix}=\phi_{\gamma(x)}^{a}(x)=\begin{bmatrix}R_{\theta^{\prime}}&0\\ 0&R_{\theta^{\prime}}\end{bmatrix}\begin{bmatrix}y\\ z\\ h_{y}\\ h_{z}\end{bmatrix}+\begin{bmatrix}y^{\prime}\\ z^{\prime}\\ 0\\ 0\end{bmatrix}.$
This gives the solution
We can then compute the group inverse of $\gamma(x)$ as:
Substituting $\gamma(x)$ into the definition of $\rho$ from (2), we get:
II-B Dimension Reduction
Using the map $\rho$ from elements in $\mathcal{X}$ to their canonical forms, we show that the dynamics in (1) are equivalent to a function operating on a lower-dimensional input space.
@@ II-B Dimension Reduction
$F:\mathcal{X}\times\mathcal{U}\to\mathcal{X}$ in system (1) is $G$-invariant if and only if there exists $\bar{F}:\mathcal{X}^{b}\times\mathcal{U}\to\mathcal{X}$ such that
@@
Assume that the dynamics $F$ are $G$-invariant. Note that $\rho$ restricted to $\mathcal{C}$ is invertible. Define $\bar{F}$ by $\bar{F}(\bar{x},u)=F(\rho|_{\mathcal{C}}^{-1}(\bar{x}),u)$. Letting $x=\rho|_{\mathcal{C}}^{-1}(\bar{x})$, this can equivalently be written as $\bar{F}(\rho(x),u)=F(x,u)$ when $x\in\mathcal{C}$. Since $F$ is $G$-invariant, $\phi_{\gamma(x)}(F(x,u))=F(\phi_{\gamma(x)}(x),\psi_{\gamma(x)}(u))$. Note that $\phi_{\gamma(x)}(x)\in\mathcal{C}$. This can be rewritten in terms of $\bar{F}$ as $\bar{F}(\rho(\phi_{\gamma(x)}(x)),\psi_{\gamma(x)}(u))$. By Lemma 1, this equals $\bar{F}(\rho(x),\psi_{\gamma(x)}(u))$, as desired.
To prove the reverse direction, assume $\bar{F}:\mathcal{X}^{b}\times\mathcal{U}\to\mathcal{X}$ exists such that (7) holds for all $x\in\mathcal{X}$, $u\in\mathcal{U}$. Let $g\in G$. Then,
$F($ $\phi_{g}(x),\psi_{g}(u))$
$=\phi_{\gamma(\phi_{g}(x))}^{-1}(\bar{F}(\rho(\phi_{g}(x)),\psi_{\gamma(\phi_{g}(x))}(\psi_{g}(u))))$
$=\phi_{\gamma(\phi_{g}(x))}^{-1}(\bar{F}(\rho(x),\psi_{\gamma(\phi_{g}(x))g}(u))).$
From the proof of Lemma 1, we know $\gamma(\phi_{g}(x))g=\gamma(x)$. Therefore, the above can be simplified to
$F(\phi_{g}(x),\psi_{g}(u))$ $=\phi_{\gamma(x)g^{-1}}^{-1}(\bar{F}(\rho(x),\psi_{\gamma(x)}(u)))$
$=\phi_{g}(\phi_{\gamma(x)^{-1}}(\phi_{\gamma(x)}(F(x,u))))$
$=\phi_{g}(F(x,u)).$
@@ III Learning Model Dynamics
A critical part of the theory in Section II is that the only knowledge of system dynamics required is that of the state space, the control input space, and the system symmetries. We leverage this to learn a dynamical model that is invariant to a given transformation group using Theorem 1. Unlike a standard formulation, which would learn the model $F:\mathcal{X}\times\mathcal{U}\to\mathcal{X}$, we learn a function $\bar{F}:\mathcal{X}^{b}\times\mathcal{U}\to\mathcal{X}$ and construct $F$ from:
@@
Note that, in this section, we use $(x,u,x^{\prime})$ to denote a state, an action applied to the state, and the next state.
The relationship between $F$ and $\bar{F}$ is depicted in Figure 1. From Theorem 1, $F$ is, by construction, $G$-invariant. A side-benefit is that training is done with a lower dimensional input space ($\mathcal{X}^{b}\times\mathcal{U}$ instead of $\mathcal{X}\times\mathcal{U}$). To learn this function $\bar{F}$, given a dataset $\mathcal{D}$ consisting of tuples $(x,u,x^{\prime})$, we train $\bar{F}$ to map $(\rho(x),\psi_{\gamma(x)}(u))$ to $\phi_{\gamma(x)}(x^{\prime})$.
@@
Note that, in practice, it is typical to learn a function which maps $(x,u)$ to $x^{\prime}-x$ instead of $x^{\prime}$ directly. To do this, we learn $\Delta\bar{F}$ to map $(\rho(x),\psi_{\gamma(x)}(u))$ to $\phi_{\gamma(x)}(x^{\prime})-\phi_{\gamma(x)}(x)$. Then, we construct $F$ by
@@
This construction for $F$ is $G$-invariant since it is equivalent to using $\bar{F}(\bar{x},\bar{u})\triangleq\Delta\bar{F}(\bar{x},\bar{u})+\rho|_{\mathcal{C}}^{-1}(\bar{x})$, which satisfies the condition in Theorem 1. Finally, $\Delta F(x,u)$ can be defined as $\Delta F(x,u)\triangleq F(x,u)-x$ to provide the same $x^{\prime}-x$ interface expected by many libraries. This construction for $\Delta F$ in terms of $\Delta\bar{F}$ expands as follows.
@@
One special case where this simplifies is when $\phi_{g}$ is a homomorphism with respect to $+$ for all $g\in G$. Then, (9) simplifies to
@@
An example where this occurs is when $\phi_{g}$ is linear.